three times a first number decreased by a second number is 22the first number increased by twice the second number is 19
Answer by ewatrrr(24785)

 
Hi,
Let x be the first number and y represent the second number:
Writing as we read:
3x - y = 22
x + 2y = 19
Solving by elimination:
6x - 2y = 44 |mutiplying first EQ by 2 to eliminate y
x + 2y = 19
7x = 63
x = 9 |3x - y = 22 OR 3x-22 = y , y = 3*9 - 22 = 5